How many different teams of 5 can be formed from a group of 8 bowlers?
Answer by edjones(8007)

8C5=8!/[(8-5)!5! Combination of 8 things taken 5 at a time.
=8!/3!*5!
=8*7*6*5!/3!*5!
=8*7*6/3*2
=8*7
=56
.
Ed